# Comparison of Probabilistic Structural Reliability Methods for Ultimate Limit State Assessment of Wind Turbines

### v1

@@ 4.2.4 Simulation of response from GP-model
Given the GP-model that enables drawing random samples of the distribution parameters for any long-term parameter $\bm{x}$, a "full" Monte-Carlo estimate of the response distribution is obtained. First, distribution parameters $\boldsymbol{\theta}_{j}$ are sampled from the GP-model for each long-term condition $x_{j}$, $j=1,\cdots,10000\cdot 365.25\cdot 24\cdot 6$ in the 10 000-year long-term simulation. Then, for each $\boldsymbol{\theta}_{j}$ a short-term response $y_{j}$ is sampled from the gumbel- og GEV distribution $\hat{g}_{Y|\bm{X}}(y|\boldsymbol{\theta}_{j})$. From the 10 000 years of responses, the relevant return values are estimated.
@@ 5 Summary and conclusions
The second case study – South Brittany – only included one contour method (IFORM), and also did not include brute force results since this was too computationally costly (essentially, because the cut-off values in the long-term parameter space would be too small). However, this case study illustrates potential difficulties and shortcomings with the environmental contour method when the short-term variability is large. Results from the environmental contours and the sequential sampling do not concur; the contour estimates of the 50-year response are significantly lower than the estimates from the sequential sampling approach. Most likely, this is because the short-term variability is dominating in this case study, violating the implicit assumptions of the environmental contour method. This is confirmed by Fig. 16, which shows that for the South Brittany case, the wind conditions most likely to be responsible the long-term extreme response lie well within the environmental contour lines. Hence, the assumption that the largest response will occur in the most severe wind conditions is not true. Indeed, due to large short-term variability, significant contributions to the long-term extreme response come from frequently occurring non-extreme wind conditions. This observation is also substantiated by the observation that there are relatively small differences between the 1-year and the 50-year extreme response estimates.

### v2

@@ 4.2.4 Simulation of response from GP-model
Given the GP-model that enables drawing random samples of the distribution parameters for any long-term parameter $\bm{x}$, a "full" Monte-Carlo estimate of the response distribution is obtained. First, distribution parameters $\boldsymbol{\theta}_{j}$ are sampled from the GP-model for each long-term condition $x_{j}$, $j=1,\cdots,10000\cdot 365.25\cdot 24\cdot 6$ in the 10 000-year long-term simulation. Then, for each $\boldsymbol{\theta}_{j}$ a short-term response $y_{j}$ is sampled from the Gumbel- or GEV distribution $\hat{g}_{Y|\bm{X}}(y|\boldsymbol{\theta}_{j})$. From the 10 000 years of responses, the relevant return values are estimated.
@@ 5 Summary and conclusions
The second case study – South Brittany – only included one contour method (IFORM), and also did not include brute force results since this was too computationally costly (essentially, because the cut-off values in the long-term parameter space would be too small). However, this case study illustrates potential difficulties and shortcomings with the environmental contour method when the short-term variability is large. Results from the environmental contours and the sequential sampling do not concur; the contour estimates of the 50-year response are significantly lower than the estimates from the sequential sampling approach. Most likely, this is because the short-term variability is dominating in this case study, violating the implicit assumptions of the environmental contour method. This is confirmed by Fig. 16, which shows that for the South Brittany case, the wind conditions most likely to be responsible the long-term extreme response lie well within the environmental contour lines. They correspond to recurrent operational conditions that will not be identified by contour methods. Hence, the assumption that the largest response will occur in the most severe wind conditions is not true. Indeed, due to large short-term variability, significant contributions to the long-term extreme response come from frequently occurring non-extreme wind conditions. This observation is also substantiated by the observation that there are relatively small differences between the 1-year and the 50-year extreme response estimates.
@@
One possible explanation for the relatively stronger influence of the short-term variability of the response in one of the case studies (South Brittany) can be that for this case study, a lower number of seeds was used for the response calculations for each environmental condition. If the extreme loading is dominated by extrapolation of the loads, a lower number of seeds could yield larger variability of the extremes. Hence, using 100 seeds rather than 1000 seeds is a possible explanation for the difference in short-term variability. Notwithstanding, even 100 seeds is a relatively large number and it is believed to correspond to a sufficient number short term response calculations. Further investigation into the effect of number of seeds is recommended in actual design applications.